Cost of Normal Spoilage Frieling Company installs granite countertops in customers' homes. First, the customer chooses the particular granite slab, and then Frieling measures the countertop area at the customer's home, cuts the granite to that shape, and installs it. The Tramel job calls for direct materials of $2,200 and direct labor of $1,000. Overhead is applied at the rate of 110 percent of direct labor cost. Unfortunately, one small countertop breaks during installation and Frieling must cut another piece and install it to properly complete the job. The additional rework required direct materials costing $700 and direct labor costing $400. Assume that the spoilage was due to carelessness by a Frieling worker and it is considered to be normal spoilage. Required: 1. Calculate the cost of the Tramel job. $ 2. Make any needed journal entry to the overhead control account. If an amount box does not require an entry, leave it blank. 3. What if the additional rework required $800 of direct labor? What would be the effect on the cost of the Tramel job?
Answer #1

1. cost of the Tramel job = direct material + direct labor + overhead = 2200 + 1000 + (110/100 * 1000) = 4300

2.

The amount for rework is mentioned in question and here is the journal entry

Journal entry
Overhead control 1100
Direct materials 700
Direct labor 400

3. Rework cost = direct material + direct labor + overhead = 700 + 400 + 440 = 1540

If 800 of direct labor required then cost = direct material + new direct labor + increased overhead= 700 + 800 + 880 = 2380

Increase amount = old cost - new cost = 2380 - 1540 = 840

If the additional rework required $800 of direct labor rather than $400, then the total cost of the Tramel job would increase by $840 to $2380.